a body starts from rest and with a uniform acceleration of 10 metre per second square for 5 seconds during then next 10 seconds it moves with uniform velocity the total distance travelled by a body

Grade:11

1 Answers

Kanishk
6 years ago
S=ut+half at square. U=0 S= 1/2 ×10×5×5=125Distance with acceleration 10m/sec square is 125 units then velocity v at that point =u+at. U=0 V=10×5=50units. Then from there constant velocity is maintained for 10sec v=d/t. d=vtd=50×10 =500units total distance =500+125=625 units